# Active Search for Low-altitude UAV Sensing and Communication for Users at Unknown Locations

### v1

@@ Abstract
This paper studies optimal unmanned aerial vehicle (UAV) placement to ensure line-of-sight (LOS) communication and sensing for a cluster of ground users possibly in deep shadow, while the UAV maintains backhaul connectivity with a base station (BS). The key challenges include unknown user locations, uncertain channel model parameters, and unavailable urban structure. Addressing these challenges, this paper focuses on developing an efficient online search strategy which jointly estimates channels, guides UAV positioning, and optimizes resource allocation. Analytically exploiting the geometric properties of the equipotential surface, this paper develops an LOS discovery trajectory on the equipotential surface while the closed-form search directions are determined using perturbation theory. Since the explicit expression of the equipotential surface is not available, this paper proposes to locally construct a channel model for each user in the LOS regime utilizing polynomial regression without depending on user locations or propagation distance. A class of spiral trajectories to simultaneously construct the LOS channels and search on the equipotential surface is developed. An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel. Numerical results on real 3D city maps demonstrate that the proposed scheme achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search.
@@
UAV trajectory design, unknown user locations, unknown channels, unknown environment, perturbation theory
@@ I Introduction
In dense urban environments, short-wavelength signals such as millimeter wave (mmW), Terahertz (THz), and optical waves experience significant attenuation due to obstructions like buildings and other structures [1]. This attenuation is particularly problematic for sensing performance, which depends heavily on maintaining line-of-sight (LOS) links between targets and transceivers [2, 3].
@@
Deploying a low-altitude unmanned aerial vehicle (UAV) presents a promising solution for establishing LOS links for users in deep shadowed areas, as UAV s operate at relatively high altitudes with the flexibility of 3D mobility. However, optimizing UAV deployment remains challenging due to the arbitrarily complex terrain topology, which complicates modeling and predicting LOS conditions. Previous works have circumvented this challenge by assuming distance-dependent pure LOS channel models [4, 5], or probabilistic channel models [6, 7, 8, 9]. While these simplified models allow for coarse analysis, they often compromise performance by neglecting the actual LOS conditions. Some state-of-the-art works have developed blockage-aware channel models [10, 11] to enhance UAV placement strategies. For instance, the authors in [10] approximated buildings as polyhedrons and derived explicit expressions to determine the LOS status of a 3D position using an offline city map. However, constructing geometric maps or radio maps incurs substantial costs, and polyhedral approximation, while useful, remains a coarse and computationally intensive method, making it challenging to scale to larger areas.
@@
In addition, most existing studies [4, 5, 6, 7, 8, 9, 10, 11, 12] assume perfect knowledge of user locations and channel models. However, in practical scenarios, accurate user locations or precise channel models are typically unavailable due to the location privacy of the users and the uncertain non-line-of-sight (NLOS) condition. Some recent research efforts have considered first estimating user locations and channel parameters before optimizing UAV service positions [13, 14, 15]. Nevertheless, localization under potential NLOS conditions is already a challenging problem, and the performance of UAV placement is highly dependent on the accuracy of user localization. Assuming the availability of 3D maps, the work in [15] employed the Fisher information metric to design an online trajectory for joint path-loss parameter estimation and user localization, although the global optimality and complexity of the approach remain unclear. With known user locations and maps, the authors in [13] utilized dynamic programming to optimize UAV trajectory for channel parameter learning, employing map compression to estimate LOS probability and assist in UAV placement. However, the effectiveness of the placement heavily depends on channel learning and LOS probability estimation. In the absence of prior knowledge of channels and user locations, a model-free approach presented in [14] relied on signal-to-interference-and-noise ratio (SINR) measurements and a 3D map to account for blockages and scatterings, applying Q-learning to optimize UAV placement. Despite this, the action space was limited to four orthogonal horizontal directions, significantly reducing the likelihood of identifying the optimal UAV placement in a 3D environment.
@@
This paper investigates the UAV trajectory design to provide communication and sensing services for a cluster of ground users at unknown locations, while also maintaining backhaul connectivity with a remote base station (BS). The goal is to develop an optimization framework for an efficient online search of the best UAV position for a broad class of communication problems and sensing problems. The challenges arise from the need to operate without precise information: user locations are unknown due to privacy concerns and potential NLOS conditions, the channel model parameters are difficult to determine due to environmental sensitivity, and the lack of accurate 3D maps further complicates the UAV placement in dense urban environments.
@@
In this work, we develop three key techniques to address the aforementioned challenges. First, we introduce a search method on an equipotential surface using trajectories derived from perturbation theory. The equipotential surface represents a region where the sensing and communication performance for users is balanced with the capacity of the UAV-BS backhaul link. The UAV’s proximity to the equipotential surface can be assessed by measuring the signal-to-noise ratio (SNR), without requiring knowledge of user locations or channel parameters. Although the exact expression of the equipotential surface remains elusive, perturbation theory guides the UAV search direction while maintaining its position on this surface. Second, we propose a local construction of channel models for each user within the LOS regime using local polynomial regression. This approach allows for a relatively simple construction of local LOS models without requiring user location or propagation distance. Third, we develop a search strategy on the equipotential surface that leverages two universal properties: upward invariance and colinear invariance of LOS regions over almost any terrain structure. Prior studies have shown that, in certain cases, this strategy can achieve an $\epsilon$-optimal solution globally in 3D space with a trajectory length proportional to the search radius [16, 17]. However, previous work [17] assumed knowledge of user locations and channel model parameters.
Contributions of this paper are as follows:
  • We study UAV trajectory design for serving a cluster of users at unknown locations with unknown channels. Specifically, we develop our search strategy based on perturbation theory and two universal properties, i.e., upward invariance and colinear invariance, of the LOS regions over almost all terrain structure.
@@
  • We provide an optimal design for a class of UAV search trajectories to balance the estimation for user channels and the exploration for better channel quality. This also includes deriving the upper bound of the mean squared error (MSE) for the locally constructed channel gain based on an optimal measurement pattern and determining the optimal radius of the spiral trajectory for minimizing the MSE of channel gain estimation.
@@
  • We demonstrate through numerical results that the normalized channel gain construction error is on the order of $10^{-2}$ without knowing user locations or propagation distances. Furthermore, with a $3$ kilometer search, the proposed scheme can achieve over $95\%$ performance of that from a 3D exhaustive search over a dense urban area.
@@
The remaining part of the paper is organized as follows: Section II introduces the system model and formulates a UAV-assisted sensing and communication problem. Section III discusses the geometric properties of the equipotential surface and local channel model construction. Section IV outlines the LOS discovery trajectory and proposes a superposed trajectory for optimal LOS position search. Section V presents numerical results and comparisons, while Section VI concludes the paper.
@@ II-A Environment Model
Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{3}:x_{3}\geq H_{\text{min}}\}$ be the feasible region of the UAV positions. While signals of ground users are likely blocked by buildings, we denote $\mathcal{D}_{k}\text{$\subseteq$}\mathcal{X}$ as the region of UAV positions such that there is an LOS link between the UAV at position $\mathbf{x}\in\mathcal{D}_{k}$ and user $k$ at an unknown position $\mathbf{u}_{k}$.
@@ II-C UAV-assisted Sensing and Communication with Backhaul
where $f_{0}(g_{0}(\mathbf{x}))$ represents the objective of the BS-UAV link under the LOS condition, $F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ represents the objective of the UAV-user links under LOS, and $H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0$ are the corresponding constraints for the resource allocation. Problem 5 is non-convex due to the LOS constraint on UAV positions and the fact that the blockage can have an arbitrary shape.
@@ II-C1 Balancing problem
Consider to deploy a UAV to offer sensing and communication services for a cluster of sensing targets and communication users while establishing a LOS relay link with a remote BS. Specify $p_{k}$ as the power allocation from the UAV to user $k$. Denote $f_{k}(g_{k}(\mathbf{x}),p_{k})$ as the objective function for user $k$, and $N_{0}$ as the noise power of the propagation channel. For sake of brevity, let $N_{0}=1$ in this paper. For a sensing task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as weighted SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{s}}p_{k}g_{k}(\mathbf{x})$ where $\mu_{\text{s}}$ is a weight of the sensing task [23, 24]. For a communication task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as channel capacity, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{c}}\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$ where $\mu_{\text{c}}$ is a weight of the communication task. In addition, the objective function of the BS-UAV link is given by the capacity function, i.e., $f_{0}(g_{0}(\mathbf{x}))=\log_{2}(1+P_{0}g_{0}(\mathbf{x}))$ where $P_{0}$ is the transmit power of the BS. The balancing problem aims at maximizing the worst link performance of the BS-UAV link and the UAV-user links. Hence, the overall objective function is specified as $\min\{f_{0}(g_{0}(\mathbf{x})),\min_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k})\}\}$. The balancing problem jointly optimizes the UAV position $\mathbf{x}$ and the power allocation $\mathbf{p}$ as follows
@@ II-C2 Sum-rate problem
While the balancing problem is non-convex and the sum-rate problem is convex in the resource allocation variable $\mathbf{p}$, the whole problem $\mathscr{P}$ that jointly optimizes the UAV location $\mathbf{x}$ is non-convex. In addition, as the channels and the LOS regions $\mathcal{D}_{k}$ are not available before exploring near location $\mathbf{x}$, the UAV needs to design an online trajectory to explore the LOS opportunity, measure the channel quality, and optimize for the system performance.
@@ III-B Trajectory towards the Equipotential Surface
III-C Construction of a Local Channel Model
It is well-known that directly estimating a global propagation model $g_{k}(\mathbf{x})$ is difficult, as it leads to nonlinear regression. Instead, for each node $k$, we adopt a linear model $\hat{g}(\mathbf{x})$ to locally approximate the LOS channel gain $g_{k}(\mathbf{x})$ for the LOS region $\mathbf{x}\in\mathcal{D}_{k}$ in the neighborhood of $\mathbf{x}=\mathbf{c}_{0}$:
@@ III-E Measurement Trajectory Design
where $\omega=4k\pi/M$ and $v=2/\sqrt{M^{2}-1}$, with $k$ being a natural number, typically $k=1$ (see Fig. 2 (a)). One can easily verify that if we sample at $t=m-1/2$, i.e., $\mathbf{x}_{m}=\mathbf{x}(m-1/2)$, for $m=1,2,\dots,M$, then conditions (i)–(iii) in Theorem 1 are satisfied. As a result, the MSE of the locally reconstructed channel $\hat{g}_{k}(\mathbf{x})$ is upper bound by (19).
@@ IV-C2 Implementation
$\dot{\mathbf{x}}_{\text{t}}=\frac{\tau-t+t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})+\frac{t-t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+}),\quad t\in(t_{1},t_{1}+\tau)$ (43)
@@
A sample implementation is summarized in Algorithm 1. As Algorithm 1 is an online search scheme, the following two metrics are essential for understanding its complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course. Specifically, given a common realization of channel models and a specified objective function for the balancing problem in Proposition 2, the trajectory length $L$ is found to be linear in some geometry parameters of the environment, i.e.,
@@
where $H_{0}$ is the initial search altitude, $R$ is the radius of $\mathcal{S}$ given in (11), and $r_{1}$ is the measurement range defined in Section III. Recall that $K$ is the number of users, $M$ is the number of measurements used for local channel model construction, and $N$ is the number of constraints for resource allocation. The per-step computational complexity is found as $O(KM+(K+N)^{3})$, which is numerically found to be handled in milliseconds. For detailed complexity derivations, see Appendix E: Supplementary Material 1.
@@
    Local channel construction: Collect channel measurements for each time slot $\triangle t$. Construct the local channel $\hat{g}_{k}(\mathbf{x})$ and obtain the estimate for $\nabla g_{k}(\mathbf{x})$ for each user based on model (16) with parameter (17) estimated from the past $M$ LOS measurements.

### v4

@@ Abstract
This paper studies optimal unmanned aerial vehicle (UAV) placement to ensure line-of-sight (LOS) communication and sensing for a cluster of ground users possibly in deep shadow, while the UAV maintains backhaul connectivity with a base station (BS). The key challenges include unknown user locations, uncertain channel model parameters, and unavailable urban structure. Addressing these challenges, this paper focuses on developing an efficient online search strategy which jointly estimates channels, guides UAV positioning, and optimizes resource allocation. Analytically exploiting the geometric properties of the equipotential surface, this paper develops an LOS discovery trajectory on the equipotential surface while the closed-form search directions are determined using perturbation theory. Since the explicit expression of the equipotential surface is not available, this paper proposes to locally construct a channel model for each user in the LOS regime utilizing polynomial regression without depending on user locations or propagation distance. A class of spiral trajectories to simultaneously construct the LOS channels and search on the equipotential surface is developed. An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel. Numerical results on real 3D city maps demonstrate that the proposed scheme achieves over $94\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search.
@@
UAV placement, unknown user locations, perturbation theory, trajectory design, channel map construction
@@ I Introduction
In dense urban environments, short-wavelength signals like millimeter wave (mmW), Terahertz (THz), and optical waves suffer from severe attenuation due to signal blockage from buildings and other structures [1]. In particular, sensing performance heavily relies on line-of-sight (LOS) links between targets and transceivers, which are usually disrupted in urban environments [2, 3].
@@
Deploying a low-altitude unmanned aerial vehicle (UAV) is a promising solution for providing LOS links for users in deep shadow since the UAV operates at a relatively high altitude with 3D mobility. Nevertheless, the optimization of UAV deployment poses challenges due to the arbitrarily complex terrain topology, which complicates modeling and predicting LOS conditions. Previous works bypassed this challenge by assuming distance-dependent pure LOS channel models [4, 5], or probabilistic channel models [6, 7, 8, 9]. These simplified channel models enable a coarse analysis but sacrifice the performance by ignoring the actual LOS condition. Some state-of-the-art works formulated blockage-aware channel models [10, 11] to assist UAV placement. For example, the authors in [10] approximated buildings as polyhedrons, and derived explicit expressions to determine the LOS status of a 3D position using an offline city map. However, the performance of these methods highly relies on the accuracy of the model of the environment.
@@
In addition, most existing works [4, 5, 6, 7, 8, 9, 10, 11, 12] assume perfect knowledge of user locations and channel models. Nevertheless, in practical scenarios, accurate user locations and precise channel models are typically unavailable due to the location privacy of users and the uncertain non-line-of-sight (NLOS) conditions. Some recent works considered first estimating user locations and channel parameters before optimizing UAV service positions [13, 14, 15]. However, localization itself under possibly NLOS conditions is already a challenging problem and the performance of UAV placement highly depends on the localization accuracy of the users. Assuming the availability of 3D maps, the work [15] employed the fisher information metric to design an online trajectory for joint path-loss parameter estimation and user localization, but the global optimality and complexity remain unclear. With known user locations and maps, the work [13] utilized dynamic programming to optimize UAV trajectory for channel parameters learning and employed map compression to estimate the LOS probability for UAV positions. However, the placement performance relies on channel learning and LOS probability estimation. Without the prior knowledge of channels and user locations, a model-free approach in [14] relied on signal-to-interference-and-noise ratio (SINR) measurements and a 3D map to account for blockage and scatters, and the authors applied Q-learning to optimize UAV placement. However, the action space is restricted to four orthogonal horizontal directions on a 2D plane, and there is little theoretical guarantee on the performance.
@@
This paper investigates the UAV trajectory design to provide communication and sensing services for a cluster of ground users at unknown locations, while the UAV maintains backhaul connectivity with a remote base station (BS). The goal is to provide an optimization framework for an efficient online search of the best UAV position for a broad class of communication problems and sensing problems. The challenges arise from the need to operate without precise information, where the user locations and the channel models are not available at the UAV, and the UAV has to make search decisions and adjust the search trajectory on-the-fly based on the online measurements.
@@
In this work, we develop three techniques to resolve the above challenges. First, we propose to search on an equipotential surface using trajectories developed from perturbation theory. The equipotential surface is the region of UAV positions where the sensing and communication performance for the users is balanced with capacity of the UAV-BS backhaul link. The status whether the UAV is on or off the equipotential surface can be quantified by measuring the signal-to-noise ratio (SNR), without requiring knowledge of user locations or channel parameters. While an explicit expression of the equipotential surface is not available, we employ perturbation theory to develop a search strategy so that the search trajectory remains on the equipotential surface. Second, we propose to locally construct a channel map for each user within the LOS regime using local polynomial regression. This approach allows for a relatively simple construction of nonparametric channel models that are capable of capturing actual signal attenuations due to blockage from the possibly unknown propagation environment. Third, we develop our search strategy on the equipotential surface exploiting two universal properties: upward invariance and colinear invariance of LOS regions over almost all terrain structure. Prior studies have shown that, in certain cases, this strategy can achieve an $\epsilon$-optimal solution globally in 3D space with a trajectory length linear of the search radius [16, 17]. Yet, the prior work [16, 17] assumed knowledge of user locations and channel model parameters.
The contributions of this paper are summarized as follows:
  • We analytically show that the equipotential surface is a sphere for a class of sensing and communication problems, where the sphere parameters depend on the user distribution and the power budget.
@@
  • We develop a class of spiral trajectories to simultaneously construct the local LOS channels and search on the equipotential surface. An optimal radius of the spiral and an optimal measurement pattern for channel gain estimation are derived to minimize the mean squared error (MSE) of the locally constructed channel.
@@
  • We demonstrate that the normalized channel gain construction error is on the order of $10^{-2}$ without knowing the user location or the propagation distance. With a $3$-kilometer search, the proposed scheme can achieve over $94\%$ performance of that from a 3D exhaustive search for a UAV-assisted multiuser sensing and communication problem over a dense urban area.
@@
The remaining part of the paper is organized as follows: Section II introduces the system model and formulates a UAV-assisted sensing and communication problem. Section III discusses the geometric properties of the equipotential surface and local channel map construction. Section IV outlines the LOS discovery trajectory and proposes a superposed trajectory for optimal LOS position search. Section V presents numerical results and comparisons, while Section VI concludes the paper.
@@ II-A Environment Model
Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{3}:x_{3}\geq H_{\text{min}}\}$ be the feasible region of UAV positions where $H_{\text{min}}$ is the minimum flight height. While signals of ground users are likely blocked by buildings, we denote $\mathcal{D}_{k}\text{$\subseteq$}\mathcal{X}$ as the region of UAV positions such that there is an LOS link between the UAV at position $\mathbf{x}\in\mathcal{D}_{k}$ and user $k$ at an unknown position $\mathbf{u}_{k}$.
@@ II-C UAV-assisted Sensing and Communication with Backhaul
where $f_{0}(g_{0}(\mathbf{x}))$ represents the objective of the BS-UAV link under the LOS condition, $F_{\text{u}}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})$ represents the objective of the UAV-user links under LOS, and $H_{n}(\bm{g}_{\text{u}}(\mathbf{x}),\mathbf{p})\leq 0$ for $n=1,2,\dots,N$ are the corresponding constraints for the resource allocation. Problem $\mathscr{P}$ is non-convex due to the LOS constraint on UAV positions and the fact that the blockage can have an arbitrary shape.
@@ II-C1 Balancing problem
Consider to deploy a UAV to offer sensing and/or communication services for a cluster of sensing targets and/or communication users while establishing a LOS relay link with a remote BS. Specify $p_{k}$ as the power allocation from the UAV to user $k$. Denote $f_{k}(g_{k}(\mathbf{x}),p_{k})$ as the objective function for user $k$, and $N_{0}$ as the noise power of the propagation channel. For sake of brevity, let $N_{0}=1$ in this paper. For a sensing task involving estimation, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as weighted SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{s}}p_{k}g_{k}(\mathbf{x})$ where $\mu_{\text{s}}$ is a weight of the sensing task [23, 24]. For a communication task, $f_{k}(g_{k}(\mathbf{x}),p_{k})$ can be specified as channel capacity, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=\mu_{\text{c}}\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$ where $\mu_{\text{c}}$ is a weight of the communication task. In addition, the objective function of the BS-UAV link is given by the capacity function, i.e., $f_{0}(g_{0}(\mathbf{x}))=\log_{2}(1+P_{0}g_{0}(\mathbf{x}))$ where $P_{0}$ is the transmit power of the BS. The balancing problem aims at maximizing the worst link performance of the BS-UAV link and the UAV-user links. Hence, the overall objective function is specified as $\min\{f_{0}(g_{0}(\mathbf{x})),\min_{k\in\mathcal{K}}\{f_{k}(g_{k}(\mathbf{x}),p_{k})\}\}$. The balancing problem jointly optimizes the UAV position $\mathbf{x}$ and the power allocation $\mathbf{p}$ as follows
@@ II-C2 Sum-rate problem
While the example formulation (6) is non-convex in the power allocation variable $\mathbf{p}$ and the problem (7) is convex in $\mathbf{p}$, both cases can be handled in a same way in our proposed algorithm framework. In addition, as the channels and the LOS regions $\mathcal{D}_{k}$ are not available before exploring near location $\mathbf{x}$, the UAV needs to design an online trajectory to explore the LOS opportunity, measure the channel quality, and optimize for the system performance.
@@ III-B Trajectory towards the Equipotential Surface
III-C Construction of a Local Channel Map
It is well-known that directly estimating a global propagation model $g_{k}(\mathbf{x})$ is difficult, as it leads to nonlinear regression. Instead, for each node $k$, we adopt a linear model $\hat{g}(\mathbf{x})$ to locally approximate the LOS channel map $g_{k}(\mathbf{x})$ for the LOS region $\mathcal{D}_{k}$ in the neighborhood of $\mathbf{x}=\mathbf{c}_{0}$:
@@ III-E Measurement Trajectory Design
where $\omega=4k\pi/M$ and $v=2/\sqrt{M^{2}-1}$, with $k$ being a natural number, typically $k=1$ (see Fig. 2 (a)). One can easily verify that if we sample at $t=m-1/2$, i.e., $\mathbf{x}_{m}=\mathbf{x}(m-1/2)$, for $m=1,2,\dots,M$, then conditions (i)–(iii) in Theorem 1 are satisfied. As a result, the MSE of the locally reconstructed channel $\hat{g}_{k}(\mathbf{x})$ is upper bounded by (19).
@@ IV-C2 Implementation
$\dot{\mathbf{x}}_{\text{t}}=\frac{\tau-t+t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{-})+\frac{t-t_{1}}{\tau}\dot{\mathbf{x}}_{\text{s}}(t_{1}^{+}),\>t\in(t_{1},t_{1}+\tau)$ (43)
@@
A sample implementation is summarized in Algorithm 1.
IV-C3 Complexity Analysis
As Algorithm 1 is an online search scheme, we investigate two different metrics for an understanding of its complexity: the trajectory length of the UAV and the per-step computational complexity every time the UAV adjusts its course.
The following proposition shows the trajectory length is linear in the radius of the equipotential surface under certain conditions.
Proposition 3 (Upper bound of trajectory length).
Suppose that $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$. For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the trajectory length $L$ is upper bounded as
@@ IV-C3 Complexity Analysis
where $H_{0}$ is the initial search altitude, $R$ is the radius of the equipotential surface given in (11), and $r_{1}$ is the measurement range.
Proof.
See Appendix E in [27]. ∎
Proposition 3 suggests that reducing the initial search altitude decreases the upper bound of the trajectory length $L$. Additionally, the trajectory length $L$ is related to the radius $R$ of the equipotential surface. In a special case where ${\bf u}_{0}=[0,0,0]^{\text{T}}$, $K=1$, and $P_{0}>P_{\text{T}}>0$, $R^{2}$ in (11) simplifies to $R^{2}=P_{0}/(P_{0}-P_{\text{T}})\left(P_{0}/(P_{0}-P_{\text{T}})-1\right)d^{2}({\bf u}_{0},{\bf u}_{1})$, which indicates that $R$ is linear in $d({\bf u}_{0},{\bf u}_{1})$. Thus, the worst-case search trajectory length is linear in the BS-user distance $d({\bf u}_{0},{\bf u}_{1})$.
Recall that $K$ is the number of users, $M$ is the number of measurements used for local channel construction, and $N$ is the number of constraints for resource allocation in problem (22). The computational complexity of Steps 3 and 4 in Algorithm 1 is found as $O(KM+(K+N)^{3})$. See Appendix F in [27] for a detailed derivation of the computational complexity.
@@
    Local channel construction: Collect channel measurements for each time slot $\triangle t$. Construct the local channel $\hat{g}_{k}(\mathbf{x})$ and obtain the estimation of $\nabla g_{k}(\mathbf{x})$ for each user based on model (16) with parameters (17) estimated from the past $M$ LOS measurements.